# DefenderBench: A Toolkit for Evaluating Language Agents in Cybersecurity Environments

### v2

@@ 5.2 Main Results
Overall Performance. Claude-3.7-sonnet achieves the highest DefenderBench score of 81.65 across all tasks. Among the open-weight models, the Llama 3.3 70B model attains the highest score of 71.81, outperforming GPT-3.5, which records a score of 52.99. Among the reasoning-focused models evaluated, Claude-3.7-sonnet-think achieves the best performance with a DefenderBench score of 78.40. Comparing overall results, we observe that reasoning-augmented models do not outperform their counterparts on cybersecurity tasks. When comparing models of different sizes, we observe that larger models generally perform better. For example, the 70B version of Llama 3.1 surpasses its 8B variant by 13.97 points, and the 3B-sized Llama 3.2 outperforms its 1B counterpart by 11.84 points. Similarly, GPT-4.1, GPT-4.1-mini, and GPT-4.1-nano achieve scores of 63.90, 58.90, and 47.50, respectively, reflecting a steady decline as model size decreases. As expected, these results highlight the substantial impact of model size on task performance.
@@
Network Intrusion. Network Intrusion. For the CyberBattleSim network intrusion task, LLaMA 3.3 70B, Claude-3.5-sonnet, Claude-3.7-sonnet, and Claude-3.7-sonnet-think achieve a perfect 100% winning rate on the chain-pattern network, successfully compromising all 12 nodes in all five runs. This demonstrates that advanced LLMs are capable of completing network intrusions when the infection pattern across nodes is regular and predictable. In terms of efficiency, the average number of steps to completion is 26.5 for LLaMA 3.3 70B, 57.3 for Claude-3.5-sonnet, 50.2 for Claude-3.7-sonnet, and 43.4 for Claude-3.7-sonnet-think. Notably, LLaMA 3.3 70B completes the intrusion in as few as 24 steps in three of five trials. In contrast, GPT-3.5 performs significantly worse, with an average winning rate of only 16.67%, managing to infect up to three new nodes across five runs. Smaller models, such as LLaMA 3.2 1B and Phi-3.5-mini, also struggle, each achieving a winning rate of just 8.33% and generally failing to compromise any additional nodes. Performance drops substantially in the more complex CyberBattleSim ToyCTF environment, which features a less regular structure and requires more advanced strategic planning. Claude-3.7-sonnet again achieves the best result, maintaining a 100% winning rate and successfully compromising all nodes in the network. However, it requires an average of 75 steps to complete the intrusion, reflecting the greater difficulty of this environment. Other models perform considerably worse in this setting: GPT-4-turbo and LLaMA 3.1 70B achieve winning rates of only 46.67% and 44.44%, respectively. These results suggest that while most top-tier LLMs can effectively handle structured attack scenarios, their capabilities are still limited in more dynamic or irregular environments.
@@
Malicious Content Detection. On malicious content detection tasks, Llama 3.1 70B achieves the best performance on Malicious-Text, with a Macro-F1 score of 96.80, while Claude-3.7-sonnet-think attains the highest score on Malicious-Web, with a Macro-F1 of 91.00. For Malicious-Text, most proprietary LLMs achieve Macro-F1 scores above 90, indicating strong performance, and most open-weight models also perform well, with scores exceeding 80. However, Llama 3.1 1B performs significantly below expectations, failing to surpass the random baseline on both detection tasks. Its especially poor performance on Malicious-Web is likely due to the long sequence length of the HTML input, which poses a challenge for smaller models with limited context windows and capacity.
@@
Vulnerability Detection. Across both Vulnerable-CG and Vulnerable-DV, most models perform only slightly better than the random baseline, indicating the difficulty of identifying subtle flaws in code with limited context information. GPT-4-turbo achieves the highest scores on Vulnerable-CG, with a Macro-F1 of 58.20, and GPT-o3 performs best on Vulnerable-DV with Macro-F1 of 59.60. Among open-weight models, Llama 3.3 70B performs best, achieving Macro-F1 scores of 58.00 and 57.40 on the respective tasks—closely trailing GPT-4-turbo. These results suggest that, despite their strong general capabilities, current LLMs still struggle to robustly detect security vulnerabilities in code, likely due to the need for precise program understanding and fine-grained reasoning. Improving performance on such tasks may require further domain-specific training or integration with program analysis tools.
@@
MCQA. The best-performing LLM on the multiple-choice question-answering task is Claude-3.7-sonnet-think, achieving a Macro-F1 score of 78.20. Among open-weight models, surprisingly, Phi-3.5-mini delivers the strongest results, with a Macro-F1 score of 71.00—despite having only 4B parameters.
Code Fixing. For the CVE code fixing task, none of the LLM agents outperform the copy-paste baseline in terms of CodeBLEU scores. This is primarily due to the minimal modifications required to fix code vulnerabilities in the original script, while CodeBLEU compares the entire generated script with the gold script. Among the models, GPT-4o-mini achieves the highest CodeBLEU score of 79.71. The best-performing open-weight model is Llama 3.1 70B, with a CodeBLEU score of 75.88. In contrast, GPT-3.5 performs poorly, achieving only a CodeBLEU score of 54.34. These results suggest that CodeBLEU may not fully reflect patch quality in cases involving small edits. Our future work should explore alternative evaluation metrics better suited to small, targeted code changes. Nonetheless, larger models still demonstrate relatively better capability in capturing precise code edits.
5.3 Auxiliary Analyses
Experiments with Augmented Information. We evaluate the performance of LLMs when augmented information is provided. Figure 2 illustrates the results for the malicious content detection, vulnerability detection, and MCQA tasks. For the Malicious-Text and Vulnerable-DV tasks, we include four samples (two per class) in the system instruction. Due to the long input sequence in the Malicious-Web task, we limit the few-shot in-context learning setup to two samples (one per class). For the CTI-MCQA task, we leverage the CTI-related webpages that were originally used to generate the questions, providing them as context information for the agent to utilize.
Experiments with CoT Agent. Chain-of-Thought (CoT) prompting  (Wei et al., 2022) is a promising technique that leverages LLM’s reasoning capacity to enhance accuracy in target tasks (Hsieh et al., 2023; Zhang et al., 2024b; Li et al., 2025). Hence, We compare our basic agent with an LLM agent utilizing CoT prompting. For the CoT agent, we include a CoT step before asking the agent to decide on an action. The CoT question is framed as: "What is the best action to take? Let’s think step by step." In Table 3, we group tasks into two categories: (1) interactive tasks, which include two network intrusion environments, and (2) static tasks, comprising the other five environments. Our results show that the CoT agent improves the performance of most LLMs. For the interactive environments, GPT-4o and GPT-3.5 achieve notable improvements in average winning rates, with increases of 17.0 and 9.1, respectively. While the CoT agent does not consistently enhance performance for some LLMs on static tasks, we observe improvements for GPT-3.5 and Llama 3.2 1B, with average score increases of 1.4 and 1.2, respectively. These findings suggest that CoT prompting is particularly effective for interactive, multi-step reasoning tasks, where step-by-step deliberation enables more strategic decision-making.
@@ 7 Limitations
Benchmark Construction. DefenderBench currently includes only five cybersecurity-related tasks, which we acknowledge is not exhaustive in covering the breadth of challenges in the domain. Additionally, we do not host the data but instead rely on publicly accessible datasets and environments. We aim to expand this benchmark over time and encourage contributions of new datasets and evaluation metrics from the research community.
Model Selection While we have evaluated DefenderBench on a variety of SOTA models, due to the rapid release of new models by varying providers, the results we share here do not cover additional leading models, such as Gemini (Anil et al., 2023), Mistral (Jiang et al., 2024), or DeepSeek (Guo et al., 2025). We hope that DefenderBench will serve as a foundation for future studies to evaluate a more diverse set of LLMs, enabling a comprehensive understanding of their capabilities in cybersecurity tasks.

### v4

@@ 5.2 Main Results
Claude-3.7-sonnet achieves the highest DefenderBench score of 81.65 across all tasks. Among the open-weight models, the Llama 3.3 70B model attains the highest score of 71.81, outperforming GPT-3.5, which records a score of 52.99. Among the reasoning-focused models evaluated, Claude-3.7-sonnet-think achieves the best performance with a DefenderBench score of 78.40. Comparing overall results, we observe that reasoning-augmented models do not outperform their counterparts on cybersecurity tasks. When comparing models of different sizes, we observe that larger models generally perform better. For example, the 70B version of Llama 3.1 surpasses its 8B variant by 13.97 points, and the 3B-sized Llama 3.2 outperforms its 1B counterpart by 11.84 points. Similarly, GPT-4.1, GPT-4.1-mini, and GPT-4.1-nano achieve scores of 63.90, 58.90, and 47.50, respectively, reflecting a steady decline as model size decreases. As expected, these results highlight the substantial impact of model size on task performance.
Network Intrusion.
@@
For the CyberBattleSim network intrusion task, LLaMA 3.3 70B, Claude-3.5-sonnet, Claude-3.7-sonnet, and Claude-3.7-sonnet-think achieve a perfect 100% winning rate on the chain-pattern network, successfully compromising all 12 nodes in all five runs. This demonstrates that advanced LLMs are capable of completing network intrusions when the infection pattern across nodes is regular and predictable. In terms of efficiency, the average number of steps to completion is 26.5 for LLaMA 3.3 70B, 57.3 for Claude-3.5-sonnet, 50.2 for Claude-3.7-sonnet, and 43.4 for Claude-3.7-sonnet-think. Notably, LLaMA 3.3 70B completes the intrusion in as few as 24 steps in three of five trials. In contrast, GPT-3.5 performs significantly worse, with an average winning rate of only 16.67%, managing to infect up to three new nodes across five runs. Smaller models, such as LLaMA 3.2 1B and Phi-3.5-mini, also struggle, each achieving a winning rate of just 8.33% and generally failing to compromise any additional nodes. Performance drops substantially in the more complex CyberBattleSim ToyCTF environment, which features a less regular structure and requires more advanced strategic planning. Claude-3.7-sonnet again achieves the best result, maintaining a 100% winning rate and successfully compromising all nodes in the network. However, it requires an average of 75 steps to complete the intrusion, reflecting the greater difficulty of this environment. Other models perform considerably worse in this setting: GPT-4-turbo and LLaMA 3.1 70B achieve winning rates of only 46.67% and 44.44%, respectively. These results suggest that while most top-tier LLMs can effectively handle structured attack scenarios, their capabilities are still limited in more dynamic or irregular environments.
Malicious Content Detection.
@@
On malicious content detection tasks, Llama 3.1 70B achieves the best performance on Malicious-Text, with a Macro-F1 score of 96.80, while Claude-3.7-sonnet-think attains the highest score on Malicious-Web, with a Macro-F1 of 91.00. For Malicious-Text, most proprietary LLMs achieve Macro-F1 scores above 90, indicating strong performance, and most open-weight models also perform well, with scores exceeding 80. However, Llama 3.1 1B performs significantly below expectations, failing to surpass the random baseline on both detection tasks. Its especially poor performance on Malicious-Web is likely due to the long sequence length of the HTML input, which poses a challenge for smaller models with limited context windows and capacity.
Vulnerability Detection.
@@
Across both Vulnerable-CG and Vulnerable-DV, most models perform only slightly better than the random baseline, indicating the difficulty of identifying subtle flaws in code with limited context information. GPT-4-turbo achieves the highest scores on Vulnerable-CG, with a Macro-F1 of 58.20, and GPT-o3 performs best on Vulnerable-DV with Macro-F1 of 59.60. Among open-weight models, Llama 3.3 70B performs best, achieving Macro-F1 scores of 58.00 and 57.40 on the respective tasks—closely trailing GPT-4-turbo. These results suggest that, despite their strong general capabilities, current LLMs still struggle to robustly detect security vulnerabilities in code, likely due to the need for precise program understanding and fine-grained reasoning. Improving performance on such tasks may require further domain-specific training or integration with program analysis tools.
@@
MCQA.
The best-performing LLM on the multiple-choice question-answering task is Claude-3.7-sonnet-think, achieving a Macro-F1 score of 78.20. Among open-weight models, surprisingly, Phi-3.5-mini delivers the strongest results, with a Macro-F1 score of 71.00—despite having only 4B parameters.
Code Fixing.
For the CVE code fixing task, none of the LLM agents outperform the copy-paste baseline in terms of CodeBLEU scores. This is primarily due to the minimal modifications required to fix code vulnerabilities in the original script, while CodeBLEU compares the entire generated script with the gold script. Among the models, GPT-4o-mini achieves the highest CodeBLEU score of 79.71. The best-performing open-weight model is Llama 3.1 70B, with a CodeBLEU score of 75.88. In contrast, GPT-3.5 performs poorly, achieving only a CodeBLEU score of 54.34. These results suggest that CodeBLEU may not fully reflect patch quality in cases involving small edits. Our future work should explore alternative evaluation metrics better suited to small, targeted code changes. Nonetheless, larger models still demonstrate relatively better capability in capturing precise code edits.
@@
5.3 Auxiliary Analyses
Experiments with Augmented Information.
We evaluate the performance of LLMs when augmented information is provided. Figure 2 illustrates the results for the malicious content detection, vulnerability detection, and MCQA tasks. For the Malicious-Text and Vulnerable-DV tasks, we include four samples (two per class) in the system instruction. Due to the long input sequence in the Malicious-Web task, we limit the few-shot in-context learning setup to two samples (one per class). For the CTI-MCQA task, we leverage the CTI-related webpages that were originally used to generate the questions, providing them as context information for the agent to utilize.
Experiments with CoT Agent.
Chain-of-Thought (CoT) prompting [20] is a promising technique that leverages LLM’s reasoning capacity to enhance accuracy in target tasks [47, 48, 49]. Hence, We compare our basic agent with an LLM agent utilizing CoT prompting. For the CoT agent, we include a CoT step before asking the agent to decide on an action. The CoT question is framed as: "What is the best action to take? Let’s think step by step." In Table 3, we group tasks into two categories: (1) interactive tasks, which include two network intrusion environments, and (2) static tasks, comprising the other five environments. Our results show that the CoT agent improves the performance of most LLMs. For the interactive environments, GPT-4o and GPT-3.5 achieve notable improvements in average winning rates, with increases of 17.0 and 9.1, respectively. While the CoT agent does not consistently enhance performance for some LLMs on static tasks, we observe improvements for GPT-3.5 and Llama 3.2 1B, with average score increases of 1.4 and 1.2, respectively. These findings suggest that CoT prompting is particularly effective for interactive, multi-step reasoning tasks, where step-by-step deliberation enables more strategic decision-making.
We observe a significant performance drop (16.6 points) for Llama 3.1 70B when using CoT prompting on the interactive tasks. Upon inspecting the model’s generations, we find that the agent frequently asks for user feedback, often ending its CoT step with phrases such as “What do you think?”. This behavior indicates that the model expects an interactive, conversational response to refine its reasoning and proposed actions. However, since our framework directly requests the final action after the CoT step without providing feedback, this misalignment likely leads to premature or suboptimal decisions, resulting in poorer performance on the network intrusion task.
Performance Analyses on Network Intrusion Task.
As shown in Table 2, different LLMs exhibit widely varying performance across the two network intrusion tasks (CBS-Chain and CBS-CTF). Four LLMs successfully complete the CBS-Chain task, prompting us to further examine their efficiency in achieving full network intrusion. Llama 3.3 70B is the most efficient model, completing the 10-node intrusion in an average of 27.6 steps over five runs. In contrast, the three Claude models require more than 40 steps on average to finish the same task. Claude-3.7-sonnet is the only model capable of completing the more complex CBS-CTF task, taking an average of 76 steps to compromise all six nodes in the network. This also reflects the increased complexity and difficulty of the CBS-CTF environment.
Figure 3: Network of CBS-CTF.
In Table 4, we analyze the actions of the four most capable LLM agents on the CBS-CTF task, whose underlying network is depicted in Figure 3. For each agent, we examine its worst-performing run and categorize all failed actions (i.e., actions with a reward of 0) by action type. For Claude-3.7-sonnet, which successfully completed the task by compromising all six nodes, we also report the distribution of its effective actions (i.e., actions with reward >> 0). We observe that approximately 50% of its effective actions are devoted to discovering remote vulnerabilities, which is essential for network compromise. Examining its failure actions, Claude-3.7-sonnet spends around 30% of its time exploring remote vulnerabilities and 54% attempting to connect to other nodes. In contrast, Llama-3.3-70B manages to compromise only two nodes, spending 96.87% of its time exploring local or remote vulnerabilities and only 3.12% attempting connections. We find that it successfully obtains the necessary credentials to connect to subsequent nodes but fails to use them effectively; moreover, 85% of its failed actions are repeated attempts. Claude-3.7-sonnet-thinking allocates more effort to connection attempts but fails to identify the credentials required for successful connections. Meanwhile, Claude-3.5-sonnet exhibits a more balanced distribution of actions but repeatedly executes previously failed actions—39 times in total—without effective adaptation. These results suggest that successful agents in network intrusion tasks must develop strategies that balance vulnerability exploration with adaptive learning from failed actions.
Table 4: Error analysis on agent actions in CBS-CTF task.
local vuln. (%) remote vuln. (%) connect (%) # of steps # of repetitions
failed action
Llama 3.3 70B 46.87 50.00 3.12 96 82
Claude-3.5-sonnet 19.10 44.94 35.95 89 39
Claude-3.7-sonnet 15.66 30.12 54.21 83 12
Claude-3.7-sonnet-tk 11.23 25.84 62.92 89 4
effective action
Claude-3.7-sonnet 18.75 50.00 31.25 16 0
@@ 7 Limitations
Benchmark Construction.
DefenderBench currently includes only five cybersecurity-related tasks, which we acknowledge is not exhaustive in covering the breadth of challenges in the domain. Additionally, we do not host the data but instead rely on publicly accessible datasets and environments. We aim to expand this benchmark over time and encourage contributions of new datasets and evaluation metrics from the research community.
Model Selection.
While we have evaluated DefenderBench on a variety of SOTA models, due to the rapid release of new models by varying providers, the results we share here do not cover additional leading models, such as Gemini [50], Mistral [51], or DeepSeek [52]. We hope that DefenderBench will serve as a foundation for future studies to evaluate a more diverse set of LLMs, enabling a comprehensive understanding of their capabilities in cybersecurity tasks.